2. Set up an integral to find the volume of the solid generated when the region bounded by y = 2x2 and y = x3 is

(a) Rotated about the x-axis using shells

(b) Rotated about the x-axis using washers

(c) Rotated about the y-axis using shells

(d) Rotated about the y-axis using washers

(e) Rotated about the line x = −3

(f) Rotated about the line y = −2

(g) Rotated about the line y = 11

(h) Rotated about the line x = 7
Answer #1

Set X3 2x c2,) 2 sing Sholls For thisArsa wi funtion f y (a) Retated about x-akis (2,8) gs Ay= 2y) (2.-2n (2-yb) 2 3/2 (Rta(x3-2x) AC 2T X Ad V= 2x3) d 2 rt uirng uasher bRotated about -xs ) AC) 7t 4)dx V cd Rtated aboct y-anis using washurs Ag)E7Rated abprtneime geit2 (Retatd about the line y -2 utey adius- -2-2x 9 1adios = -2-x3 M2x A) 2-)-2-22 8 -2K3 C9 Rotated about

Rotatedabot ne 2x2 gs +3 Inner ndius 33 ,(28) 3 3+3, 2 Vs h) Rotated about line 7 Outer 91adius Mner sadiuss7 (2,9 VE 7